OpenStudy (anonymous):

What is 27x2 − 13x2?

OpenStudy (anonymous):

2*14=28

OpenStudy (anonymous):

that is not a answer

OpenStudy (anonymous):

\[(27*2)-(13*x)=2(27-13)=28\]

OpenStudy (anonymous):

54-26=28

OpenStudy (anonymous):

u mean x^2?

OpenStudy (anonymous):

is 54 - 26 = 28. This is due to operator precedence. Remember BODMAS (B)rackets (O)rders (powers, square roots) (DM) Division, multiplication (AS) Addition, subtraction in that order

OpenStudy (anonymous):

(27-13)x^2 14x^2

OpenStudy (anonymous):

x is a variable in tehre or its bump??

OpenStudy (anonymous):

54-26=28